OpenStudy (anonymous):

A box contains four letter tiles with the letters M, A, T, and H. A spinner has five equal sections numbered 1 through 5. What is the probability of drawing a tile labeled T from the box and spinning an odd number on the spinner?

OpenStudy (anonymous):

A.3/20 B.1/3 C.4/9 D.17/20

OpenStudy (anonymous):

@what!? @ryamorgan284 @tylermcmullen23 @YungKhris @uri @iamabarbiegirl @oliviadetataaa

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thanks

OpenStudy (welshfella):

thge probability of getting a T on first draw = 1/4 because there are 4 letters in the box. there are 3 odd numbers out of 5 on the spinner so probability of getting an odd is 3/5

OpenStudy (welshfella):

these 2 events are independent so you multiply the required probability = 1/4 * 3/5 = 3 / 20

OpenStudy (welshfella):

A is the correct answer

OpenStudy (anonymous):

can u help with one more @welshfella

OpenStudy (welshfella):

yes but please post a new question